OpenStudy (anonymous):

y/24-2=x what is the range?

OpenStudy (anonymous):

(y/24)-2=x ===> y/24 = x+2 ===> y =(1/24)x-(1/12) ===> range is Real Number Set

OpenStudy (anonymous):

(y/24)−2=x−−−>(y/24=x+2)−−−>Y=(24x+48) R:(−∞,+∞)

OpenStudy (anonymous):

the range in the book is all eral numbers except for -1 and -2

OpenStudy (anonymous):

\[\frac{y}{24} - 2 = x\] \[y = 24(x+2)\]  Since x can assume any value, y also can assume any value for a certain x, so the range is the real set. \[Range = IR\]

OpenStudy (anonymous):

now give me a good answer medal!

OpenStudy (anonymous):

the rang and domain of such function whuch first degree function and represents a stright line is alway all Real Number Set